OpenStudy (anonymous):

solve and check the inequality, if possible [4m+8]<12

OpenStudy (anonymous):

I need an answer A.S.A.P

OpenStudy (anonymous):

Well, you want to isolate the variable. So subtract 8 from each side to get 4m < 4. Now you divide each side by 4 to get m < 1. To check this, think of a number less than 1. Just take 0 to be simple. 4*0 + 8 = 0 + 8 = 8, which IS less than 12. So there's your answer :)

OpenStudy (anonymous):

stay on and I will do the problem and then let me know if I got it

myininaya (myininaya):

Are those brackets around 4m+8 Does it have any special meaning?

OpenStudy (anonymous):



OpenStudy (anonymous):

Check this way: your solution is m < 1. So take any number less than one and plug it into the equation in place of m. If the resulting inequality is true then you have the right answer.

OpenStudy (anonymous):

it's for a test and i can't even make 1+2=3 when they want it as 2+1=3 or I'll get points taken so just let me know how to check my answer and i'll be good to go